Workers: Simon, Miyo, Akutsu
- Continued to calibrate the photosensors.
With different current sources to the LEDs; OSEM satellite box and Agilent F3630A; the calibration factors seemed to change just according to the supplied current; see here
About the issue yesterday, we found the difference of the calibration factor shapes were not due to the difference of the reflectors surface; see here
- Determined our mind to install the photosensors anyhow.
- Locked the suspension.
Installed the heat-link anchor block and all the heat links, finding nice/good alignment of the heat links... see here
[Ongoing] Need cares of water springs from anchor bolts for oplev pylons.
[Ongoing] Need to confirm what are the scratch lines made by the measurement by the KEK team last week.
- Complete the calibration of the oplev; the arm length was reported by Simon, so....
- Check the balance of the suspended WAB with some wires and sensors are attached.
- Check how the damping would be now after the replacement of the damping magnets.
- Check the reproducibility of the alignment of the suspension among several modes.
- Check heights of several points.
- Take care of water spring??
- Two burndy connectors, vacuum compatible.
Stages outside the chamber to ease wiping our legs.
Good rail(?) to move the WAB suspension in the cryostat in the main beam direction.
Chain block? to support the load in the chamber.
- A clock to know the time easily.
- M6 screw set.